# Stability-Certified On-Policy Data-Driven LQR via Recursive Learning and Policy Gradient

### v1

@@ 2.2.2 Model-based gradient method for problem (5)
The set of stabilizing gains $\mathcal{K}$ is open [46, Lemma IV.3] and connected [46, Lemma IV.6], therefore, if the pair $({A}_{\star},{B}_{\star})$ were known, the gradient descent method could be used to solve problem (5) (see, e.g., [30]). Namely, at each iteration $t\in\mathbb{N}$, an estimate $K_{t}$ of $K^{\star}$ is maintained and iteratively updated according to
@@ 3 On-policy LQR for Unknown Systems: Concurrent Learning and Optimization
Initialization: $\mathrm{x}_{0}\in\mathbb{R}^{n}$, $H_{0}\in\mathbb{R}^{(n+m)\times(n+m)}$, $S_{0}\in\mathbb{R}^{(n+m)\times n}$, $\theta_{0}\!\in\!\mathbb{R}^{(n+m)\times n},K_{0}\!\in\!\mathbb{R}^{m\times n}$ and $\mathrm{w}_{0}\!\in\!\mathbb{R}^{n_{\mathrm{w}}}$.
@@
$H_{{t+1}}$ $=\lambda H_{t}+\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ \mathrm{u}_{t}\end{bmatrix}\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ \mathrm{u}_{t}\end{bmatrix}^{\top}$ (9e)
$S_{{t+1}}$ $=\lambda S_{t}+\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ \mathrm{u}_{t}\end{bmatrix}\mathrm{y}_{t}$ (9h)
$\theta_{{t+1}}$ $=\theta_{t}-\gamma H_{t}^{\dagger}\left(H_{t}\theta_{t}-S_{t}\right).$ (9i)
@@
Data from the controlled system (1d) are recast in an identification-oriented form described by
@@
$\theta_{{t+1}}=\theta_{t}-\gamma$ $\left(\sum_{\tau=0}^{t}\lambda^{t-\tau}\mathcal{H}(\mathrm{x}_{\tau},\mathrm{u}_{\tau})\right)^{\dagger}\times\left(\sum_{\tau=0}^{t}\lambda^{t-\tau}\left(\mathcal{H}(\mathrm{x}_{\tau},\mathrm{u}_{\tau})\theta_{t}-\mathcal{S}(\mathrm{x}_{\tau},\mathrm{u}_{\tau},\mathrm{y}_{\tau})\right)\right),$
where $\mathcal{H}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{(n+m)\times(n+m)}$ and $\mathcal{S}:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}^{(n+m)\times n}$ reads as
@@
The estimate $\theta_{t}$ is concurrently exploited in the update of the feedback gain $K_{t}$, replacing the unavailable $\theta^{\star}$ into (7) giving rise to (10). To ensure sufficiently informative data, we equip our feedback policy with an additive dithering signal $\mathrm{d}_{t}\in\mathbb{R}^{m}$. Namely, we implement
@@
where $\mathrm{w}_{t}\in\mathbb{R}^{n_{\mathrm{w}}}$, with $n_{\mathrm{w}}\geq n+m$, is the state of the exogenous system having $F\in\mathbb{R}^{n_{\mathrm{w}}\times n_{\mathrm{w}}}$ and $E\in\mathbb{R}^{n_{\mathrm{w}}}$ as state and output matrix, respectively. The matrix $F$ is a degree of freedom to properly shape the oscillation frequency of $\mathrm{w}_{t}$. The following assumption formalizes the requirements for the design of system (3).
@@
The signals $\mathrm{w}_{t}$ and $\mathrm{d}_{t}$ are persistently exciting and sufficiently rich of order $({n}+1)$, respectively, i.e., there exist $\alpha_{1},\alpha_{2},t_{\mathrm{w}},\mathbf{t}_{\mathrm{d}}>0$ such that, if $\mathrm{w}_{0}\neq 0_{n_{\mathrm{w}}}$, then
$\alpha_{1}I_{n_{\mathrm{w}}}\leq\sum_{\tau=\bar{t}+1}^{\bar{t}+t_{\mathrm{w}}}\mathrm{w}_{\tau}\mathrm{w}_{\tau}^{\top}\leq\alpha_{2}I_{n_{\mathrm{w}}},\quad\textnormal{ for all $\bar{t}\in\mathbb{N}$ }$ (18a)
$\operatorname{rank}\left(\begin{bmatrix}\mathrm{d}_{0}&\mathrm{d}_{1}&\ldots&\mathrm{d}_{\mathbf{t}_{\mathrm{d}}-n-1}\\ \mathrm{d}_{1}&\mathrm{d}_{2}&&\mathrm{d}_{\mathbf{t}_{\mathrm{d}}-n}\\ \vdots&\vdots&\ddots&\vdots\\ \mathrm{d}_{n}&\mathrm{d}_{n+1}&\ldots&\mathrm{d}_{\mathbf{t}_{\mathrm{d}}-1}\end{bmatrix}\right)=m(n+1).$ (18f)
Moreover, the eigenvalues of $F$ lie on the unit disk. normal-$\blacksquare$
We point out that recent references, see, e.g., [48, 14], refer to the property (18f) as persistency of excitation of order $(n+1)$, while we used the equivalent definition of sufficient richness of order $(n+1)$, see, e.g., [49].
@@
$H_{{t+1}}$ $=\lambda H_{t}+\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}+E\mathrm{w}_{t}\end{bmatrix}\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}^{\top}$ (19g)
$S_{{t+1}}$ $=\lambda S_{t}+\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}^{\top}\theta^{\star}$ (19l)
$\theta_{{t+1}}$ $=\theta_{t}-\gamma H_{t}^{\dagger}\left(H_{t}\theta_{t}-S_{t}\right)$ (19m)
$K_{{t+1}}$ $=K_{t}-\gamma G(K_{t},\theta_{t}),$ (19n)
in which we have used the explicit expressions for $\mathrm{y}_{t}$ (cf. (14)) and $\mathrm{u}_{t}$ (cf. (16)). Next, we provide the main result of the paper, i.e., the convergence properties of system (3). For the sake of compactness, we introduce $\mathcal{V}:=(\mathbb{R}^{n_{\mathrm{w}}}\setminus\{0\})\times\mathbb{R}^{n}\times\mathbb{R}^{(n+m)\times(n+m)}\times\mathbb{R}^{(n+m)\times n}\times\Theta_{b}\times\mathcal{K}_{b}(K^{\star})$ and the operators $\mathrm{v}_{H}:\mathbb{R}^{(n+m)^{2}\times n_{\mathrm{w}}^{2}}\times\mathbb{R}^{n}_{\mathrm{w}}\to\mathbb{R}^{(n+m)\times(n+m)}$ and $\mathrm{v}_{S}:\mathbb{R}^{(n+m)m\times n_{\mathrm{w}}^{2}}\times\mathbb{R}^{n}_{\mathrm{w}}\to\mathbb{R}^{(n+m)\times n}$ defined as
$\mathrm{v}_{H}(\Pi_{H},\mathrm{w})$ $:=\mathtt{unvec}\left(\Pi_{H}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)\right)$
$\mathrm{v}_{S}(\Pi_{S},\mathrm{w})$ $:=\mathtt{unvec}\left(\Pi_{S}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)\right),$
where $\Pi_{H}\in\mathbb{R}^{(n+m)^{2}\times n_{\mathrm{w}}^{2}}$, $\Pi_{S}\in\mathbb{R}^{(n+m)m\times n_{\mathrm{w}}^{2}}$.
@@
Consider system (3) and let Assumptions 2.1 and 3.1 hold. Then, for each initial condition $(\mathrm{w}_{0},\mathrm{x}_{0},H_{0},S_{0},\theta_{0},K_{0})\in\mathcal{V}$ such that $A_{0}+B_{0}K_{0}$ is Schur, there exist $\Pi_{\mathrm{x}}\in\mathbb{R}^{n\times n_{\mathrm{w}}}$, $\Pi_{H}\in\mathbb{R}^{(n+m)^{2}\times n_{\mathrm{w}}^{2}}$, $\Pi_{S}\in\mathbb{R}^{(n+m)m\times n_{\mathrm{w}}^{2}}$, $a_{1},a_{2},a_{3},a_{4},a_{5},a_{6},a_{7},a_{8},\bar{\gamma}>0$ such that, it holds
$\left\|\mathrm{\mathrm{x}}_{t}-\Pi_{\mathrm{x}}\mathrm{w}_{t}\right\|\leq a_{1}\left\|\mathrm{x}_{0}-\Pi_{\mathrm{x}}\mathrm{w}_{0}\right\|\exp(-a_{2}t)$ (20a)
$\left\|H_{t}\!-\!\mathrm{v}_{H}(\Pi_{H},\!\mathrm{w}_{t})\right\|\leq a_{3}\!\left\|H_{0}\!-\!\mathrm{v}_{H}(\Pi_{H},\!\mathrm{w}_{0})\right\|\!\exp(-a_{4}t)$ (20b)
$\left\|S_{t}\!-\!\mathrm{v}_{S}(\Pi_{S},\mathrm{w}_{t})\right\|\leq a_{5}\!\left\|S_{0}\!-\!\mathrm{v}_{S}(\Pi_{S},\mathrm{w}_{0})\right\|\!\exp(-a_{6}t)$ (20c)
$\left\|\begin{bmatrix}\theta_{t}-\theta^{\star}\\ K_{t}-K^{\star}\end{bmatrix}\right\|\leq a_{7}\left\|\begin{bmatrix}\theta_{0}-\theta^{\star}\\ K_{0}-K^{\star}\end{bmatrix}\right\|\exp(-a_{8}t),$ (20h)
for all $\gamma\in(0,\bar{\gamma})$. normal-$\blacksquare$
@@
The result (20a) of Theorem 3.2 ensures that the origin is an exponentially practically stable equilibrium for (19b). Indeed Theorem 3.2 allows us to choose the initial condition of the exogenous system $\mathrm{w}_{0}$ so that $\mathrm{\mathrm{x}}_{t}$ exponentially converges into the ball $\mathcal{B}_{\rho}(0_{n})$ for any desired radius $\rho>0$. More in details, since $\mathrm{w}_{t}$ evolves according to a marginally stable oscillating dynamics (cf. Assumption 3.1), it holds $\left\|\mathrm{w}_{t}\right\|=\left\|\mathrm{w}_{0}\right\|$ for all $t\in\mathbb{N}$. Thus, in order to make $\mathcal{B}_{\rho}(0_{n})$ attractive for the trajectories of (19b), it is sufficient to choose $\mathrm{w}_{0}$ such that
$\left\|\mathrm{w}_{0}\right\|\leq\dfrac{\rho}{\left\|\Pi_{\mathrm{x}}\right\|}.$
Furthermore, the result (20h) ensures that $(K^{\star},\theta^{\star})$ is exponentially stable for (19m) and (19n). Hence, we asymptotically (i) reconstruct the unknown system matrices $({A}_{\star},{B}_{\star})$ and (ii) compute the optimal gain matrix $K^{\star}$.

### v2

@@ 2.3.2 Model-based gradient method for problem (8)
The set of stabilizing gains $\mathcal{K}$ is open [47, Lemma IV.3] and connected [47, Lemma IV.6]. Moreover, the cost function $J(\cdot,\theta^{\star})$ is coercive [30, Lemma 3.7]. Had the pair $({A}_{\star},{B}_{\star})$ been known, the gradient descent method could have been used to solve problem (8) (see, e.g., [30]). Namely, at each iteration $t\in\mathbb{N}$, an estimate $K_{t}$ of $K^{\star}$ is maintained and iteratively updated according to
@@ 3 On-policy LQR for Unknown Systems: Concurrent Learning and Optimization
Initialization: $\mathrm{x}_{0}\in\mathbb{R}^{n}$, $H_{0}\in\mathbb{R}^{(n+m)\times(n+m)}$, $S_{0}\in\mathbb{R}^{(n+m)\times n}$, $\theta_{0}\!\in\!\mathbb{R}^{(n+m)\times n},K_{0}\!\in\!\mathbb{R}^{m\times n}$, and $\mathrm{w}_{0}\!\in\!\mathbb{R}^{n_{\mathrm{w}}}$.
@@
$H_{{t+1}}$ $=\lambda H_{t}+\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ \mathrm{u}_{t}\end{bmatrix}\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ \mathrm{u}_{t}\end{bmatrix}^{\top}$ (12a)
$S_{{t+1}}$ $=\lambda S_{t}+\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ \mathrm{u}_{t}\end{bmatrix}\mathrm{y}_{t}$ (12b)
$\theta_{{t+1}}$ $=\theta_{t}-\gamma H_{t}^{\dagger}\left(H_{t}\theta_{t}-S_{t}\right).$ (12c)
@@
Data from the controlled system (5b) are recast in an identification-oriented form described by
@@
$\theta_{{t+1}}$ $=\theta_{t}-\gamma\left(\sum_{\tau=0}^{t}\lambda^{t-\tau}\mathcal{H}(\mathrm{x}_{\tau},\mathrm{u}_{\tau})\right)^{\dagger}\left(\sum_{\tau=0}^{t}\lambda^{t-\tau}\left(\mathcal{H}(\mathrm{x}_{\tau},\mathrm{u}_{\tau})\theta_{t}-\mathcal{S}(\mathrm{x}_{\tau},\mathrm{u}_{\tau},\mathrm{y}_{\tau})\right)\right),$
where $\mathcal{H}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{(n+m)\times(n+m)}$ and $\mathcal{S}:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{n}\to\mathbb{R}^{(n+m)\times n}$ read as
@@
If the algorithm can be executed with $\gamma=1$ (and with a starting batch of samples ensuring the invertibility of $\sum_{\tau=0}^{t}\lambda^{t-\tau}\mathcal{H}(\mathrm{x}_{\tau},\mathrm{u}_{\tau})$ at each iteration), one may reformulate the learning process by maintaining only one additional state, rather than $H_{t}$ and $S_{t}$, see [48]. In this regard, we also note that the optimization (13) and learning (12c) parts share the same stepsize $\gamma$. By choosing two different stepsizes, we could induce a timescale separation, which is actually unnecessary because, as will become clear from the analysis, the two systems are in cascade rather than feedback interconnected.
@@
The estimate $\theta_{t}$ is concurrently exploited in the update of the feedback gain $K_{t}$, replacing the unavailable $\theta^{\star}$ into (10) giving rise to (13).
To ensure sufficiently informative data, we equip our feedback policy with an additive dithering signal $\mathrm{d}_{t}\in\mathbb{R}^{m}$. Namely, we implement
@@
where $\mathrm{w}_{t}\in\mathbb{R}^{n_{\mathrm{w}}}$, with $n_{\mathrm{w}}\geq{n}+{m}$, is the state of the exogenous system having $F\in\mathbb{R}^{n_{\mathrm{w}}\times n_{\mathrm{w}}}$ and $E\in\mathbb{R}^{{m}\times n_{\mathrm{w}}}$ as state and output matrix, respectively. The matrix $F$ is a degree of freedom to properly shape the oscillation frequency of $\mathrm{w}_{t}$. The following assumption formalizes the requirements for the design of system (17).
@@
The signal $\mathrm{w}_{t}$ is persistently exciting, while $\mathrm{d}_{t}$ is sufficiently rich of order $({n}+1)$, i.e., there exist $\alpha_{1},\alpha_{2},t_{\mathrm{w}},\mathbf{t}_{\mathrm{d}}>0$ such that, if $\mathrm{w}_{0}\neq 0_{n_{\mathrm{w}}}$, then
$\alpha_{1}I_{n_{\mathrm{w}}}\leq\sum_{\tau=\bar{t}+1}^{\bar{t}+t_{\mathrm{w}}}\mathrm{w}_{\tau}\mathrm{w}_{\tau}^{\top}\leq\alpha_{2}I_{n_{\mathrm{w}}},$ (18a)
$\operatorname{rank}\left(\begin{bmatrix}\mathrm{d}_{\bar{t}}&\mathrm{d}_{\bar{t}+1}&\ldots&\mathrm{d}_{\bar{t}+\mathbf{t}_{\mathrm{d}}-n-1}\\ \mathrm{d}_{\bar{t}+1}&\mathrm{d}_{\bar{t}+2}&&\mathrm{d}_{\bar{t}+\mathbf{t}_{\mathrm{d}}-n}\\ \vdots&\vdots&\ddots&\vdots\\ \mathrm{d}_{\bar{t}+n}&\mathrm{d}_{\bar{t}+n+1}&\ldots&\mathrm{d}_{\bar{t}+\mathbf{t}_{\mathrm{d}}-1}\end{bmatrix}\right)=m(n+1),$ (18b)
for all $\bar{t}\in\mathbb{N}$. Moreover, the eigenvalues of $F$ in (17) lie on the unit circle. $\blacksquare$
We point out that recent references, see, e.g., [50, 14], refer to the property (18b) as persistency of excitation of order $(n+1)$, while we used the equivalent definition of sufficient richness of order $(n+1)$, see, e.g., [51].
Remark 3.2
A possible way to build $F$ and $E$ to verify Assumption 3.1 is as follows. First, we set $q:=n+1$, $n_{\mathrm{w}}=2q$, and $F:=\operatorname{blkdiag}(F_{1},\dots,F_{q}),$ where, for all $i\in\{1,\ldots,q\}$, $F_{i}\in\mathbb{R}^{2\times 2}$ is defined as
for given $\omega_{i}$ such that $\omega_{i}=2\omega_{i-1}$ for all $i\in\{2,\ldots,q\}$. By choosing an initial condition $\mathrm{w}_{0}$ that satisfies
$([\mathrm{w}_{0}]_{2i+1})^{2}+([\ w_{0}]_{2i+2})^{2}\neq 0,\quad 0\leq i\leq q-1,$
the chosen structure of $F$ guarantees (18a) according to [52, Thm. 2]. As for (18b), it can be achieved by selecting $E$ such that $\begin{bmatrix}E^{\top}&(EF)^{\top}&\ldots&(EF^{n})^{\top}\end{bmatrix}$ is nonsingular. $\blacksquare$
@@
$H_{{t+1}}$ $=\lambda H_{t}+\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}+E\mathrm{w}_{t}\end{bmatrix}\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}^{\top}$ (19c)
$S_{{t+1}}$ $=\lambda S_{t}+\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}^{\top}\theta^{\star}$ (19d)
$\theta_{{t+1}}$ $=\theta_{t}-\gamma H_{t}^{\dagger}\left(H_{t}\theta_{t}-S_{t}\right)$ (19e)
$K_{{t+1}}$ $=K_{t}-\gamma G(K_{t},\theta_{t}),$ (19f)
in which we have used the explicit expressions for $\mathrm{y}_{t}$ (cf. (14)) and $\mathrm{u}_{t}$ (cf. (16)). Next, we provide the main result of the paper, i.e., the convergence properties of system (19). For the sake of compactness, we introduce the sets $\mathcal{V}:=(\mathbb{R}^{n_{\mathrm{w}}}\setminus\{0\})\times\mathbb{R}^{n}\times\mathbb{R}^{(n+m)\times(n+m)}\times\mathbb{R}^{(n+m)\times n}\times\Theta_{b}\times\mathcal{K}_{b}(K^{\star})$ and
$\mathcal{V}_{\text{ss}}(\Pi_{1},\Pi_{2},\Pi_{3}):=\{(\mathrm{w},\mathrm{x},H,S,\theta,K)\in\mathcal{V}\mid\mathrm{x}$ $=\Pi_{1}\mathrm{w},H=\mathrm{v}_{H}(\Pi_{2}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)),$
$S$ $=\mathrm{v}_{S}(\Pi_{3}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)),\theta=\theta^{\star},$
$K$ $=K^{\star}\},$
in which we also introduced the operator $\mathrm{v}_{H}:\mathbb{R}^{(n+m)^{2}\times n_{\mathrm{w}}^{2}}\times\mathbb{R}^{n_{\mathrm{w}}}\to\mathbb{R}^{(n+m)\times(n+m)}$ defined as
$\mathrm{v}_{H}(\Pi_{1},\mathrm{w})$ $:=\mathtt{unvec}\left(\Pi_{1}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)\right),$ (20a)
and the operator $\mathrm{v}_{S}:\mathbb{R}^{(n+m)m\times n_{\mathrm{w}}^{2}}\times\mathbb{R}^{n_{\mathrm{w}}}\to\mathbb{R}^{(n+m)\times n}$ defined as
$\mathrm{v}_{S}(\Pi_{2},\mathrm{w})$ $:=\mathtt{unvec}\left(\Pi_{2}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)\right).$ (20b)
@@
Consider system (19) and let Assumptions 2.7 and 3.1 hold. Then, for each initial condition $(\mathrm{w}_{0},\mathrm{x}_{0},H_{0},S_{0},\theta_{0},K_{0})\in\mathcal{V}$ such that $A_{0}+B_{0}K_{0}$ is Schur, there exist $\Pi_{\mathrm{x}}\in\mathbb{R}^{n\times n_{\mathrm{w}}}$, $\Pi_{H}\in\mathbb{R}^{(n+m)^{2}\times n_{\mathrm{w}}^{2}}$, $\Pi_{S}\in\mathbb{R}^{(n+m)m\times n_{\mathrm{w}}^{2}}$, and $\bar{\gamma}>0$ such that, for all $\gamma\in(0,\bar{\gamma})$, the set $\mathcal{V}_{\text{ss}}(\Pi_{\mathrm{x}},\Pi_{H},\Pi_{S})$ is exponentially stable for system (19). $\blacksquare$
@@
In addition to the robustness properties intrinsically inherited with the stability, it is important to note that Theorem 3.3 also ensures exponential convergence of the states $(\mathrm{\mathrm{x}}_{t},\theta_{t},K_{t})$ toward $(\Pi_{\mathrm{x}}\mathrm{w}_{t},\theta^{\star},K^{\star})$, namely, along the trajectories of (19), it holds
$\left\|\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\!-\!\Pi_{\mathrm{x}}\mathrm{w}_{t}\\ \theta_{t}\!-\!\theta^{\star}\\ K_{t}\!-\!K^{\star}\end{bmatrix}\right\|\leq a_{1}\exp(-a_{2}t),$ (21)
for some $a_{1},a_{2}>0$. Therefore, with a linear rate, Relearn LQR asymptotically reconstructs the unknown system matrices $({A}_{\star},{B}_{\star})$ and achieves the optimal gain $K^{\star}$. We recall that $\mathrm{w}_{t}$ evolves according to a marginally stable oscillating dynamics (cf. Assumption 3.1) and, thus, $\left\|\mathrm{w}_{t}\right\|=\left\|\mathrm{w}_{0}\right\|$ for all $t\in\mathbb{N}$. Consequently, given $\rho>\left\|\Pi_{\mathrm{x}}\mathrm{w}_{0}\right\|$, result (21) ensures that the ball $\mathcal{B}_{\rho}(0_{n})$ is exponentially attractive for (19b), provided that the other states lie in their steady-state locus. In turn, this means that we can arbitrarily reduce the radius $\rho$ by acting on $\mathrm{w}_{0}$, i.e., the origin is a practically exponentially stable equilibrium for (19b), provided that the other states lie in their steady-state locus.
Remark 3.4
We point out that Theorem 3.3 represents a local result. In fact, as for the the initial pair $(A_{0},B_{0})\in\Theta_{b}$, it must belong to the largest ball centered at $({A}_{\star},{B}_{\star})$ of controllable matrices $(A,B)$. Also, an initial controller $K_{0}$ that stabilizes both $(A_{0},B_{0})$ and $({A}_{\star},{B}_{\star})$ is required. The proof of Theorem 3.3 then exploits system theory tools based on averaging theory for two-time-scale systems (cf. Theorem 2.6 in Section 2.1) and, therefore, introduces an auxiliary system called the averaged system. Such auxiliary system involves modified averaged dynamics of $\theta_{t}$ and $K_{t}$ (see Section 4.2) and is shown to have an exponentially stable equilibrium in its origin. Moreover, we are also able to show that $K_{t}$ remains a stabilizing gain for $({A}_{\star},{B}_{\star})$ at all $t\in\mathbb{N}$ (see Appendix C). Such stabilizing property combined with closeness between trajectories of the averaged and original systems, imposed through $\gamma$ (cf., e.g., [45]), allows us to conclude that $K_{t}$, generated by (19), stabilizes $({A}_{\star},{B}_{\star})$ for all $t\in\mathbb{N}$. $\blacksquare$